The side of a square equals the width of a rectangle. The length of the rectangle is 20 meters longer than its width. The sum of the areas of the square and the rectangle is 342 square meters. Find the side of the square.
Answer by ikleyn(52818)

.
x^2 + x*(x+20) = 342.

Solve for x.

2x^2 + 20x - 342 = 0,
x^2 + 10x - 171 = 0,
(x+19)*(x-9) = 0.

The only positive solution is x = 9.

Answer. The side of the square is 9 m.